OpenStudy (anonymous):

solve this system by substitution 4x=y-1 6x-2y=-3

OpenStudy (anonymous):

4x=y-1 y=4x+1

OpenStudy (anonymous):

plug that into y of other equation

OpenStudy (anonymous):

Y or x of other equation

OpenStudy (anonymous):

y

OpenStudy (anonymous):

4x=y-1 -y = -4x -1 -y/-y = -4x/- -1/- y = 4x +1 Now substitute y in the equation below and solve for x 6x-2y=-3 6x-2(4x + 1)=-3 Solve for x in the equation 6x-2(4x + 1)=-3 Once you have the value of x, put the value of x into the equation 4x=y-1 and then solve for y

OpenStudy (anonymous):

If you need anymore help let us know

OpenStudy (anonymous):

Also here is a web site if you just want the answers http://www.wolframalpha.com/input/?i=4x%3Dy-1+and+6x-2y%3D-3